# Factorisation conditions and causality for local measurements in QFT

### v1

@@ V.1 No impossible operations
In the second line above we have used the identity Eq. (57) to split each Kraus operator into those localised before and after the Cauchy slice. In the third line above we have used the localisation properties to commute $C$ with $\Pi[f_{+}]$ and $\Phi_{\textsc{a}}$ with $\Pi[f_{-}]$ (see figure4). The fourth line follows from the fact that the argument of $\Phi_{\textsc{a}}$, integrated over $x$, is proportional to the identity i.e. using Eq (58) which is shown in the appendix VII.5. Importantly, this leads to $\Phi_{\textsc{a}}$ dropping out from the expression, and as a result the details of A’s preparation disappear from C’s result.
@@ V.2 Effective Kraus operators
Finally, we use that the map $f\mapsto\Pi[f]$ is strongly continuous [oeckl2024spectraldecompositionfieldoperators]171717technically, this was proven only for the map between solutions and Kraus operators with $\psi$ Gaussian (see discussion of [oeckl2024spectraldecompositionfieldoperators] in section IV.2)., to see that observables localised to R are updated by the ‘effective’ would-be operators
@@
Note that the suggested notion of effective operation, relies on covariance, i.e. freedom to split on any spacelike Cauchy slice, and on being able to take the limit that we sketched above. This is to guarantee that, in a sense, there are no superluminal processes within the support of $f$. Importantly, resolving Sorkin’s problem does not rely on this limiting procedure, but on the fact that there is always a choice of hypersurface such that $f_{-}$ will be fully spacelike to A, and so B cannot enable a superuminal signal from A to C (see figure 4). After all, even in the case when continuous additivity is not available, impossible operations are excluded by the Hammerstein property, as shown in the appendix VII.1.
@@ V.3 Generality of the results
for $H^{\pm}$ localised in $\operatorname{supp}{f}\cap J^{\pm}(\Sigma)$. This is expected as generically only polynomials of first order in the fields lead to no-signalling unitary kicks. More details on this will be provided in [simmons_brukner_in_prep].
@@ V.4 Measuring observables beyond $\phi(f)$
We have shown that there exists a very simple solvable model that can measure $\phi(f)$ for any $f$, in a completely causal way, and which recovers the results of [PhysRevD.105.025003, albertini2023ideal, oeckl2024spectraldecompositionfieldoperators], specifically that the Kraus operators should be given by a suitable $L^{2}$-normalised function of $\phi(f)$ (in the spectral calculus sense), which in our case depends on the choice of the state of the probe. However, $\phi(f)$ is a special operator in the algebra: it is unique in that products of smeared fields will generally not generate non-signalling unitary kicks. Since the non-signalling of a Gaussian measurement of $\phi(f)$ can be shown to follow from the non-signalling of unitary kicks by $e^{i\phi(f)}$ [PhysRevD.105.025003], it would be surprising if Gaussian (and more general $L^{2}$-normalised) measurements of powers of smeared fields were non-signalling. Indeed, it has been shown for a large class of products of smeared fields, that Gaussian measurements are signalling [PhysRevD.105.025003, albertini2023ideal].
@@
Compact operators of this form can be measured in a non-signalling way [Borsten:2019cpc] with an update map that projects onto its eigenstate i.e. a non-destructive measurement. However, it is well known from quantum information [PhysRevLett.90.010402, PhysRevA.49.4331] that compact ‘non-local’ operators cannot be measured in a way that does not erase all local information. Non-local here refers to the fact that they cannot be written in a way that is completely local with respect to the tensor product decomposition e.g. $A\otimes B+A^{\prime}\otimes B^{\prime}\neq\tilde{A}\otimes 1+1\otimes\tilde{B}$. In field theory,
@@
This leads us to conclude that there are two distinct consequences of causality for measurements in QFT: 1) dynamical assumptions preclude the approximations used in the usual von Neumann derivation of ideal measurements, 2) non-local operators must be measured in such a way that the selective state update is destructive [PhysRevA.49.4331, PhysRevLett.90.010402]. While a projective measurement or even the measurement model we have studied here has Kraus operators that are simple spectral functions of the observable, and so final state is closely related to the initial state, for non-local observables, we expect the selective version of any causal update rule will lead to a final state that is very different from the original, destroying the initial state.
@@ VI Conclusions
The impossibility of importing many ‘standard’ channels into QFT leaves a large gap in its foundations. Inspired by this, we have studied the relationship between non-signalling channels and unitaries using the local $S$-matrix approach, and related it to non-signalling measurement channels in QFT. Within this approach there are natural conditions, namely a hierarchy of factorisation conditions, which are sufficient for ruling out Sorkin’s impossible operations. Further, we have studied the operational implications for the resulting channels and Kraus operators, leading to causality conditions and the notion of effective Kraus operators. This provides a bridge between the abstract channels proposed in [PhysRevD.105.025003, albertini2023ideal, oeckl2024spectraldecompositionfieldoperators] and the rigorous FV framework [fewster2020quantum, PhysRevD.103.025017, fewster_asymptotic_2023].
@@
While the $S$-matrix framework we highlight is very general, the application to even the simplest model, of a trivially evolving pointer, has been fruitful. As well as providing a purification of abstract channels, it naturally includes dynamical information that does not appear in those channels constructed by hand. Specifically, we see a lower bound on the sharpness of a measurement of $\phi(f)$ given by $\Delta_{R}(f,f)^{\frac{1}{2}}$, and thus explicit dependence on the causal connectivity of the coupling region. To reach a sharp measurement, one must replace $f\mapsto f\lambda$, and take $\lambda\rightarrow 0$, amounting to changing the physical coupling (see e.g. [mandrysch2024quantumfieldmeasurementsfewsterverch]), and while also taking the probe state to be more sharp in $Q$. Similarly, we find that our model exhibits a generalisation of the convolution identity for Kraus operators found in [oeckl2024spectraldecompositionfieldoperators]. We’ve argued that this identity is a special case of a more general property of Kraus operators, and follows from the quantum controlled nature of the $S$-matrix in the simple model.
@@
The connection between factorisation (continuous additivity and Hammerstein) and ruling out impossible measurements is not surprising, as they both address notions of causality. However, it does provide a direct link between an operational notion of no signalling in Sorkin scenarios, and a QFT notion of causality, i.e. factorisation of the $S$-matrix across any Cauchy surface. This stands in contrast with the conditions proposed in [sorkin1993impossible, Borsten:2019cpc, PhysRevD.105.025003, albertini2023ideal, Fuksa], which are much more global, in the sense that they depend non-trivially on the whole measurement scenario, and do not necessarily reduce to a condition that must hold locally within a given region, independently of the global scenario. Since the stronger form of continuous additivity implies microcausality of the Lagrangian, and microcausality of spacelike Heisenberg fields, we see a restoration of the connection between microcausality and no-signalling, which fits nicely into the AQFT framework.
@@
It is tempting at this point to comment on the possible connections to a causal Stinespring’s theorem. Of course, the Stinespring dilation theorem is a general fact about unital CP channels between $C^{*}$-algebras, and so holds in QFT abstractly. However, in the literature a causal Stinespring theorem has been eluded to several times [PhysRevD.105.025003, fewster_asymptotic_2023], well summarised by Jubb “It would be useful to construct an explicit dictionary between update maps and specific probe models, and determine if this is possible in general. The latter would be analogous to Stinespring’s Dilation theorem, but with the added restriction of locality and causality on the unitary”. Relatedly, in this paper we have shown that the abstract maps of [PhysRevD.105.025003, albertini2023ideal, oeckl2024spectraldecompositionfieldoperators], when interpreted as functionals of the function $f$, can be dilated to an $S$-matrix $S[f]$ that obeys factorisation. In this case we found that the Kraus operators have general compositional properties that are directly related to the factorisation of $S[f]$. This has the flavour of a causal Stinespring theorem, or at least can be viewed as a hint for relativistic constraints on dilated channels. Since we have only focused on a small subset of local $S$-matrices when discussing the properties of the induced Kraus operators, we cannot make general statements. However, we believe that this analysis provides hints towards a causal Stinespring theorem, something we will present general results on elsewhere [simmons_brukner_in_prep].
@@ Acknowledgements.
This research was funded in whole or in part by the Austrian Science Fund (FWF) [10.55776/F71] and [10.55776/COE1]. For open access purposes, the authors have applied a CC BY public copyright license to any author accepted manuscript version arising from this submission. This publication was made possible through the financial support of WOST (WithOutSpaceTime) grant from the John Templeton Foundation. The opinions expressed in this publication are those of the authors and do not necessarily reflect the views of the John Templeton Foundation. The authors are grateful to Robert Oeckl and José de Ramon Rivera for detailed feedback during this work. We also thank Jan Mandrysch and Miguel Navascués for discussions on the connections between our models. We are also grateful to Francisco Calderón and Doreen Fraser for insightful comments on earlier versions of this draft.
@@ VII.1 All operations with the Hammerstein property are possible
In this section we show that operations induced by local $S$-matrices obeying the first three conditions of section III.1 and the Hammerstein property [rejzner2019localitycausalityperturbativeaqft],
@@
To reach the second and third lines, we have used that $S[f_{+}+f_{0}]$ and $S[f_{0}]$ commute with $\mathfrak{A}(\text{C})$ respectively. To reach the fourth line, we have used that $S[f_{-}+f_{0}]^{\dagger}CS[f_{-}+f_{0}]$ is a product of operators that are localised spacelike to A, so $\mathcal{E}_{\textsc{a}}$ acts trivially by locality. Finally, we have used unitarity of the $S$-matrices and cyclicity of the trace to reach the final line. This implies $\delta=0$, and since we have not made any assumptions about the regions, or $\rho,C$, we can conclude that $S[f]$ is possible if it satisfies the Hammerstein property. Specifically, we can see that $\delta=0$ for all conditions implies $S[f]^{\dagger}\mathfrak{A}(\text{C})S[f],\mathfrak{A}(\text{A})]=0$ if $\mathfrak{A}(\text{C}),\mathfrak{A}(\text{A})]=0$, which implies that it is a non-signalling evolution [Borsten:2019cpc].
@@ VII.4 Non-signalling convolution property of the Kraus operators in the linear pointer coupling
where $\tilde{\psi}$ is the Fourier transform of $\psi$. This simple model satisfies the following convolution property, which was been proven to be relevant for ensuring no-signalling [oeckl2024spectraldecompositionfieldoperators]

### v2

@@ V.1 No impossible operations
In the second line above we have used the identity Eq. (57) to split each Kraus operator into those localised before and after the Cauchy slice. In the third line above we have used the localisation properties to commute $C$ with $\Pi[f_{+}]$ and $\Phi_{\textsc{a}}$ with $\Pi[f_{-}]$ (see figure4). The fourth line follows from the fact that the argument of $\Phi_{\textsc{a}}$, integrated over $x$, is proportional to the identity i.e. using Eq (58) which is shown in Appendix VII.5. Importantly, this leads to $\Phi_{\textsc{a}}$ dropping out from the expression, and as a result the details of A’s preparation disappear from C’s result.
@@ V.2 Effective Kraus operators
Finally, we use that the map $f\mapsto\Pi[f]$ is strongly continuous [35]181818technically, this was proven only for the map between solutions and Kraus operators with $\psi$ Gaussian (see discussion of [35] in section IV.2)., to see that observables localised to R are updated by the ‘effective’ would-be operators
@@
Note that the suggested notion of effective operation, relies on covariance, i.e. freedom to split on any spacelike Cauchy slice, and on being able to take the limit that we sketched above. This is to guarantee that, in a sense, there are no superluminal processes within the support of $f$. Importantly, resolving Sorkin’s problem does not rely on this limiting procedure, but on the fact that there is always a choice of hypersurface such that $f_{-}$ will be fully spacelike to A, and so B cannot enable a superuminal signal from A to C (see figure 4). After all, even in the case when continuous additivity is not available, impossible operations are excluded by the Hammerstein property, as shown in Appendix VII.1.
@@ V.3 Generality of the results
for $H^{\pm}$ localised in $\operatorname{supp}{f}\cap J^{\pm}(\Sigma)$. This is expected as generically only polynomials of first order in the fields lead to no-signalling unitary kicks. More details on this will be provided in [45].
@@ V.4 Measuring observables beyond $\phi(f)$
We have shown that there exists a very simple solvable model that can measure $\phi(f)$ for any $f$, in a completely causal way, and which recovers the results of [28, 1, 35], specifically that the Kraus operators should be given by a suitable $L^{2}$-normalised function of $\phi(f)$ (in the spectral calculus sense), which in our case depends on the choice of the state of the probe. However, $\phi(f)$ is a special operator in the algebra: it is unique in that products of smeared fields will generally not generate non-signalling unitary kicks. Since the non-signalling of a Gaussian measurement of $\phi(f)$ can be shown to follow from the non-signalling of unitary kicks by $e^{i\phi(f)}$ [28], it would be surprising if Gaussian (and more general $L^{2}$-normalised) measurements of powers of smeared fields were non-signalling. Indeed, it has been shown for a large class of products of smeared fields, that Gaussian measurements are signalling [28, 1].
@@
Compact operators of this form can be measured in a non-signalling way [6] with an update map that projects onto its eigenstate i.e. a non-destructive measurement. However, it is well known from quantum information [49, 40] that compact ‘non-local’ operators cannot be measured in a way that does not erase all local information. Non-local here refers to the fact that they cannot be written in a way that is completely local with respect to the tensor product decomposition e.g. $A\otimes B+A^{\prime}\otimes B^{\prime}\neq\tilde{A}\otimes 1+1\otimes\tilde{B}$. In field theory,
@@
This leads us to conclude that there are two distinct consequences of causality for measurements in QFT: 1) dynamical assumptions preclude the approximations used in the usual von Neumann derivation of ideal measurements, 2) non-local operators must be measured in such a way that the selective state update is destructive [40, 49]. While a projective measurement or even the measurement model we have studied here has Kraus operators that are simple spectral functions of the observable, and so final state is closely related to the initial state, for non-local observables, we expect the selective version of any causal update rule will lead to a final state that is very different from the original, destroying the initial state.
@@ VI Conclusions
The impossibility of importing many ‘standard’ channels into QFT leaves a large gap in its foundations. Inspired by this, we have studied the relationship between non-signalling channels and unitaries using the local $S$-matrix approach, and related it to non-signalling measurement channels in QFT. Within this approach there are natural conditions, namely a hierarchy of factorisation conditions, which are sufficient for ruling out Sorkin’s impossible operations. Further, we have studied the operational implications for the resulting channels and Kraus operators, leading to causality conditions and the notion of effective Kraus operators. This provides a bridge between the abstract channels proposed in [28, 1, 35] and the rigorous FV framework [16, 7, 17].
@@
While the $S$-matrix framework we highlight is very general, the application to even the simplest model, of a trivially evolving pointer, has been fruitful. As well as providing a purification of abstract channels, it naturally includes dynamical information that does not appear in those channels constructed by hand. Specifically, we see a lower bound on the sharpness of a measurement of $\phi(f)$ given by $\Delta_{R}(f,f)^{\frac{1}{2}}$, and thus explicit dependence on the causal connectivity of the coupling region. To reach a sharp measurement, one must replace $f\mapsto f\lambda$, and take $\lambda\rightarrow 0$, amounting to changing the physical coupling (see e.g. [30]), and while also taking the probe state to be more sharp in $Q$. Similarly, we find that our model exhibits a generalisation of the convolution identity for Kraus operators found in [35]. We’ve argued that this identity is a special case of a more general property of Kraus operators, and follows from the quantum controlled nature of the $S$-matrix in the simple model.
@@
The connection between factorisation (continuous additivity and Hammerstein) and ruling out impossible measurements is not surprising, as they both address notions of causality. However, it does provide a direct link between an operational notion of no signalling in Sorkin scenarios, and a QFT notion of causality, i.e. factorisation of the $S$-matrix across any Cauchy surface. This stands in contrast with the conditions proposed in [47, 6, 28, 1, 21], which are much more global, in the sense that they depend non-trivially on the whole measurement scenario, and do not necessarily reduce to a condition that must hold locally within a given region, independently of the global scenario. Since the stronger form of continuous additivity implies microcausality of the Lagrangian, and microcausality of spacelike Heisenberg fields, we see a restoration of the connection between microcausality and no-signalling, which fits nicely into the AQFT framework.
@@
It is tempting at this point to comment on the possible connections to a causal Stinespring’s theorem. Of course, the Stinespring dilation theorem is a general fact about unital CP channels between $C^{*}$-algebras, and so holds in QFT abstractly. However, in the literature a causal Stinespring theorem has been eluded to several times [28, 17], well summarised by Jubb “It would be useful to construct an explicit dictionary between update maps and specific probe models, and determine if this is possible in general. The latter would be analogous to Stinespring’s Dilation theorem, but with the added restriction of locality and causality on the unitary”. Relatedly, in this paper we have shown that the abstract maps of [28, 1, 35], when interpreted as functionals of the function $f$, can be dilated to an $S$-matrix $S[f]$ that obeys factorisation. In this case we found that the Kraus operators have general compositional properties that are directly related to the factorisation of $S[f]$. This has the flavour of a causal Stinespring theorem, or at least can be viewed as a hint for relativistic constraints on dilated channels. Since we have only focused on a small subset of local $S$-matrices when discussing the properties of the induced Kraus operators, we cannot make general statements. However, we believe that this analysis provides hints towards a causal Stinespring theorem, something we will present general results on elsewhere [45].
@@ Acknowledgements.
This research was funded in whole or in part by the Austrian Science Fund (FWF) [10.55776/F71] and [10.55776/COE1]. For open access purposes, the authors have applied a CC BY public copyright license to any author accepted manuscript version arising from this submission. This publication was made possible through the financial support of WOST (WithOutSpaceTime) grant from the John Templeton Foundation. The opinions expressed in this publication are those of the authors and do not necessarily reflect the views of the John Templeton Foundation. The authors are grateful to Robert Oeckl and José de Ramon Rivera for detailed feedback during this work. We also thank Jan Mandrysch and Miguel Navascués for discussions on the connections between our models. We are also grateful to Francisco Calderón and Doreen Fraser for insightful comments on earlier versions of this draft. Finally, the authors have benefited from the activities of COST Action CA23115: Relativistic Quantum Information, funded by COST (European Cooperation in Science and Technology) and, in particular, from discussions at EMS-IAMP Spring School: Symmetry and Measurement in Quantum Field Theory at York, United Kingdom.
@@ VII.1 All operations with the Hammerstein property are possible
In this section we show that operations induced by local $S$-matrices obeying the first three conditions of section III.1 and the Hammerstein property [42],
@@
To reach the second and third lines, we have used that $S[f_{+}+f_{0}]$ and $S[f_{0}]$ commute with $\mathfrak{A}(\text{C})$ respectively. To reach the fourth line, we have used that $S[f_{-}+f_{0}]^{\dagger}CS[f_{-}+f_{0}]$ is a product of operators that are localised spacelike to A, so $\mathcal{E}_{\textsc{a}}$ acts trivially by locality. Finally, we have used unitarity of the $S$-matrices and cyclicity of the trace to reach the final line. This implies $\delta=0$, and since we have not made any assumptions about the regions, or $\rho,C$, we can conclude that $S[f]$ is possible if it satisfies the Hammerstein property. Specifically, we can see that $\delta=0$ for all conditions implies $S[f]^{\dagger}\mathfrak{A}(\text{C})S[f],\mathfrak{A}(\text{A})]=0$ if $\mathfrak{A}(\text{C}),\mathfrak{A}(\text{A})]=0$, which implies that it is a non-signalling evolution [6].
@@ VII.4 Non-signalling convolution property of the Kraus operators in the linear pointer coupling
where $\tilde{\psi}$ is the Fourier transform of $\psi$. This simple model satisfies the following convolution property, which was been proven to be relevant for ensuring no-signalling [35]